Two cars A and B approach a road crossing with 5m per sec and 10m per sec if angle between roads is 60 degree the relative speed of car A wrt car B is

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Relative velocity of A wrt B is\vec{v_A}-\vec{v_B}

Hence relative speed= magnitude of the relative velocity=\left|\vec{v_A}-\vec{v_B}\right|

=\sqrt{v_A^2+v_B^2+2v_Av_Bcos60^{\circ}}=\sqrt{5^2+10^2+2(5)(10)\dfrac{1}{2}}=\sqrt{175}m/s